It costs Camp, Inc. $46 per unit to manufacture 1,000 units per month of a product that it can sell for $76 each. Alternatively, Camp could process the units further into a more complex product, which would cost an additional $39 per unit. Camp could sell the more complex product for $107 each. How would processing the product further affect Camp's profit?

Profit would increase by $8,000.

Profit would increase by $31,000.

Profit would decrease by $8,000.

Profit would decrease by $31,000.
Answer #1

Revenue if sold as is = (76-46)*1000 = 30,000

Revenue if sold after processing = (107 - 46 - 39)* 1000 = 22,000

So,it is clear that profits will decrease by 8,000 if processed further.

Option C is the answer